OpenStudy (anonymous):

if 14 pads of paper cost $17.50, how much will 8 pads of paper cost?

OpenStudy (anonymous):

$10

OpenStudy (anonymous):

find cost per pad, multiply by 8

OpenStudy (zzr0ck3r):

14*x = 17.50 x = 17.50/14 8*x = ?

OpenStudy (anonymous):

If the original price of the paper is 14 for $17.50 than you just divide 17.50 by 14 to find the price for each individual pad of paper. So 17.50/14 is $1.25. Meaning that each piece of paper cost $1.25. So now if you want to know how much 8 pads of paper is, you simply multiply 8 * $1.25. The answer is $10.00 :)

OpenStudy (anonymous):

If the original price of the paper is 14 for $17.50 than you just divide 17.50 by 14 to find the price for each individual pad of paper. So 17.50/14 is $1.25. Meaning that each piece of paper cost $1.25. So now if you want to know how much 8 pads of paper is, you simply multiply 8 * $1.25. The answer is $10.00 :)

OpenStudy (anonymous):

$10

OpenStudy (precal):